a liter of chemical x is 95% pure, how many liters of water should be added to make a 50% solution?
Answer by robertb(5830)

0.95*1 + 0.00*x = 0.50(1+x)
==> 0.95 = 0.50 + 0.50x
==> 0.45 = 0.50x
==> x = 0.9 liters of water should be added.